OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (-2, 4) and is parallel to y = -3x + 6.

OpenStudy (aroub):

parallel=same slope y=mx+b , m= -3 , (-2,4) 4=-3(-2)+b 4=6+b -6+4=b -2=b so the equation is y=-3x-2

OpenStudy (anonymous):

y+3x=-2